OpenStudy (anonymous):

given that 5 x -2 = -10, use a property of arithmetic to explain why it makes sense that -2 x 5 = -10.

OpenStudy (anonymous):

commutative property a*b = b*a so 5*-2 = -2*5 both are same

OpenStudy (anonymous):

commutative property of multiplication ...